### v2

@@ 3.1 ENBED outperforms state-of-the-art on genomic benchmark datasets
We finetune the model using a classification head (Sec. 2.4.2) using the embedding outputs from the final encoder block, on the datasets constructed by the authors of the Nucleotide Transformer (NT) benchmarks DallaTorre2023TheNT and Genomic Benchmarks (GB) Greov2022GenomicBA . The results of evaluating the model on the test dataset of NT and GB are presented in Tables 1 and 2, respectively. For evaluation on the NT benchmarks, we compare our performance against the Nucleotide Transformer (v2) and HyenaDNA DallaTorre2023TheNT ; Nguyen2023HyenaDNALG , which are encoder-only and decoder-only models, respectively. For the GB datasets, we use the performance of the Convolutional Neural Network (CNN) model developed by the authors of the dataset Greov2022GenomicBA as a baseline. We also include the performance of the HyenaDNA model and the baseline Transformer developed by its authors Nguyen2023HyenaDNALG .
NT Benchmark Enformer Avsec2021 DNABERT-2 zhou2023dnabert2 NT (v2) DallaTorre2023TheNT HyenaDNA Nguyen2023HyenaDNALG
@@
H3 85.9 89.3 89.5 88.9 64.4 90.6
H3K14ac 63.5 75.9 76.9 80.9 51.6 81.4
H3K36me3 67.1 79.7 81.3 80.8 61.1 82.7
H3K4me1 64.6 75.8 77.7 75.8 58.4 77.9
H3K4me2 63.0 68.0 67.6 73.9 55.9 75.7
H3K4me3 56.5 67.3 69.5 77.5 50.9 77.9
H3K79me3 74.7 80.7 81.3 83.7 83.1 85.4
H3K9ac 70.8 77.1 78.0 79.3 60.2 82.6
H4 86.6 89.9 90.5 88.2 74.3 91.8
H4ac 63.8 73.1 74.9 78.4 67.2 80.5
Promotor (all) 95.4 97.1 97.6 96.0 94.3 98.0
Promotor (non-TATA) 95.5 97.2 97.6 96.0 94.4 98.0
Promotor (TATA) 96.0 95.5 96.6 94.1 92.9 96.8
Splice acceptor 91.4 97.5 98.7 95.8 87.8 95.8
Splice donor 90.6 96.3 98.7 95.8 87.7 95.4
Enhancer 72.3 75.7 77.3 75.9 65.2 78.3
Enhancer Types 55.4 62.0 62.6 59.5 51.4 70.0
Table 1: Nucleotide Transformer (NT) Benchmarks. We evaluate our model using the 10-fold mean accuracy (%) score of the best performing variants of the Nucleotide Transformer v2 DallaTorre2023TheNT and HyenaDNA Nguyen2023HyenaDNALG , highlighting the best and second-best scores. The scores are sourced from a leaderboard maintained by the authors of DallaTorre2023TheNT on the Hugging Face platform leaderboard .
Genomic Benchmark CNN DNABERT GPT HyenaDNA Nguyen2023HyenaDNALG
@@
Mouse Enhancers 69.0 66.9 80.1 85.1 75.5 90.3
Human Enhancers (Cohn) 69.5 74.0 70.5 74.2 54.3 71.2
Human Enhancers (Ensembl) 68.9 85.7 83.5 89.2 83.3 92.2
Coding vs Intergenomic 87.6 92.5 88.8 91.3 84.2 93.0
Human vs Worm 93.0 96.5 95.6 96.6 90.8 97.3
Human Regulatory Elements 93.3 88.1 91.5 93.8 80.8 90.2
Human Promoter (Non-TATA) 84.6 85.6 87.7 96.6 83.4 97.2
Human OCR (Ensembl) 68.0 75.1 73.0 80.9 64.3 81.9
Table 2: Genomic Benchmarks. Accuracy (%) scores of the best and second-best model in the Genomic Benchmarks datasets Greov2022GenomicBA . The baseline CNN and Transformer scores was calculated by the authors of Greov2022GenomicBA and Nguyen2023HyenaDNALG respectively.
The combination of byte-level analysis, higher quality reference sequences and improved pre-training methodology helps us exceed state-of-the-art performance in 15 out of 17 of the Nucleotide Transformer (NT) benchmarks and 6 out of 8 of the Genomic Benchmarks (GB) datasets. Byte-level tokenization helps the model internalize patterns better in the presence of variations like single nucleotide polymorphisms, and the combination of encoder and decoder blocks achieves superior performance over decoder-only methods due to its ability to focus on multiple sections of the input simultaneously and to process inputs in a context-aware manner.
3.2 ENBED identifies noise in genomic sequences
@@
DNABERT jiDNABERTPretrainedBidirectional2021 84.6
Nucleotide Transformer DallaTorre2023TheNT 91.2
ENBED This paper 97.1
@@
Table 3 shows the results of the sequence-level classification on erroneous sequences using our synthetic dataset. Since competing models are trained using the GRCh38 reference assembly, they often lack information about repetitive regions due to hard-masking. We use the telomere-to-telomere assembly for pre-training the model, resulting in a signifcant performance improvement and on overall accuracy of 97.1% in the sequence-level classification task of identifying erroneous genomic data, which is significant improvement as compared to the baselines of DNABERT jiDNABERTPretrainedBidirectional2021 (84.6%) and Nucleotide Transformer DallaTorre2023TheNT (91.2%).
3.3 ENBED identifies biological function annotations
@@
DNABERT jiDNABERTPretrainedBidirectional2021 63.2
Nucleotide Transformer DallaTorre2023TheNT 67.5
HyenaDNA Nguyen2023HyenaDNALG 72.8
@@
ENBED is trained to identify the ten most common base-level annotations of the Human reference assembly. As shown in Table 4, we achieve an $F_{1}$ score of 74.1 in this classification task, an improved score compared to contemporary models DNABERT jiDNABERTPretrainedBidirectional2021 (63.2), Nucleotide Transformer DallaTorre2023TheNT (67.5), and HyenaDNA Nguyen2023HyenaDNALG (72.8).
We evaluate the accuracy of ENBED in generating mutations, using an encoder-decoder Transformer with Byte-Pair Encoding (BPE) tokenization (used in previous genomic models fishmanGENALMFamilyOpenSource2023 ) as a baseline. We compare against BPE because this method shares similarities with byte-level tokenization by starting with the basic {A, C, T, G} alphabet, but tries to optimize the vocabulary by combining simpler words into more complex ones based on the corpus the tokenizer is trained on. While this procedure reduces the average number of tokens generated from any input sequence, it also results in reduced accuracy since modifying even a single base pair will output a significantly different tokenized sequence.
Top-1 and Top-5 Accuracy (%) scores are calculated by comparing predictions with real-world data from the Influenza Virus Resource Bao2008 , with any deviation from an exact match being classified as incorrect. Top-5 scores are calculated by selecting the best candidate from the procedure described in Sec 2.5. Additionally, we also train a version of ENBED with the encoder removed, as a comparison of the sequence-to-sequence task performance between decoder-only and encoder-decoder models.
@@ 4 Ablation Studies
4.1 Encoder-decoder architecture
4.2 Architectual modifications
We also evaluate the model’s pre-training accuracy during Masked Language Modeling using dense attention and our improved linearized attention with the same GPU memory allowance, demonstrating improved performance (Table 7) due to the greater context length.
Complexity Input length Accuracy(%)
Full Attention 512 68
Linearized Attention 16384 77
Table 7: MLM Accuracy. Sub-quadratic attention outperforms full $O(N^{2})$ attention when evaluated using Masked Language Modeling (MLM) accuracy (%) with similar GPU memory usage.
5 Conclusion
We presented the ENBED foundation model, a byte-level encoder-decoder Transformer pre-trained on an ensemble of reference genomes, with a sub-quadratic attention implementation and an input length of $16,384$. We demonstrated that it generalizes well in applications involving multiple species, noisy data, and can perform a wide range of downstream tasks as summarized below:
  • Achieved state-of-the-art scores in 6 out of 8 categories on the Genomic Benchmarks datasets and 15 out of 17 categories in the Nucleotide Transformer datasets.
  • Demonstrated the ability to identify insertion, deletion and mismatch noise in sequences (Acc = 97.1%).
  • Identified biological function annotations from sequence data with an $F_{1}$ score of 74.1
  • Generated mutations in Influenza with a Top-1 and Top-5 accuracy of 76.9% and 95.4% and mean Levenshtein distance of 2.3 bp from real-world influenza mutations in 500 bp generated sequences.
6 Data and Code Availability
(GCF_000001635.27) and P. vivax (GCF_000002415.2) were obtained from NCBI RefSeq OLeary2015 in FASTA format. Variant Calling Files (VCFs) for the 1000 Genomes Project tgp were obtained from the European Bioinformatics Institute. Gene annotations were obtained from GENCODE GENCODE and Ensembl Martin2022Ensembl2 . The mutation tree was derived from the data assembled by the authors of Berman2020MutaGANAS , sourced from the NCBI’s Influenza Virus Resource Bao2008 .
The source code used to develop and fine-tune the foundation model has been released on Github. [^1] https://github.itap.purdue.edu/Clan-labs/ENBED
7 Acknowledgement
This work uses the Anvil supercomputer anvil at Purdue University through allocation CIS230228 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. This work was supported in part by the National Science Foundation under grant FW-HTF-R-2129097. The authors gratefully acknowledge the Walther Cancer Foundation and support from the Purdue University Institute for Cancer Research, P30CA023168

### v3

@@ 3.1 ENBED outperforms state-of-the-art models on genomic benchmark datasets
We finetune the model using a classification head using the embedding outputs from the final encoder block, on the datasets constructed by the authors of the Nucleotide Transformer (NT) benchmarks Dalla-Torre et al. (2023) and Genomic Benchmarks (GB) Grevsova et al. (2022). The results of evaluating the model on the test dataset of NT and GB are presented in Tables 12 and 11, respectively. For evaluation on the NT benchmarks, we compare our performance against the Nucleotide Transformer (v2) and HyenaDNA Dalla-Torre et al. (2023); Nguyen et al. (2023), which are encoder-only and decoder-only models, respectively. For the GB datasets, we use the performance of the Convolutional Neural Network (CNN) model developed by the authors of the dataset Grevsova et al. (2022) as a baseline. We also include the performance of the HyenaDNA model and the baseline Transformer developed by its authors Nguyen et al. (2023).
ENBED demonstrates superior performance, exceeding state-of-the-art results in 15 out of 17 Nucleotide Transformer (NT) benchmarks and 6 out of 8 Genomic Benchmarks (GB) datasets. This improvement likely stems from our novel approach combining byte-level analysis, high-quality reference sequences, and an optimized pre-training methodology. We hypothesize that byte-level tokenization enhances the model’s ability to handle variations such as single nucleotide polymorphisms, while our encoder-decoder architecture enables simultaneous focus on multiple input sections and context-aware processing. These features may contribute to ENBED’s advantages over decoder-only methods. While the relative impact of each component requires further investigation through ablation studies, our results demonstrate ENBED’s effectiveness across a wide range of genomic analysis tasks.
3.2 ENBED identifies noise in genomic sequences
Table 3 shows the results of the sequence-level classification on erroneous sequences using our synthetic dataset. Since competing models are trained using the GRCh38 reference assembly, they often lack information about repetitive regions due to hard-masking. Our choice of higher quality pre-training data results in a signifcant performance improvement and on overall accuracy of 97.1% in the sequence-level classification task of identifying erroneous genomic data, which is significant improvement as compared to the baselines of DNABERT Ji et al. (2021) (84.9%) and Nucleotide Transformer Dalla-Torre et al. (2023) (91.8%).
3.3 ENBED identifies biological function annotations
ENBED is trained to identify the annotations (defined in the Application Domains section) of the Human reference assembly. As shown in Table 4, we achieve an $F_{1}$ score of 74.1 in this classification task, an improved score compared to DNABERT Ji et al. (2021) (63.2), Nucleotide Transformer Dalla-Torre et al. (2023) (67.5), and HyenaDNA Nguyen et al. (2023) (72.8). For the purposes of this evaluation, all models were finetuned and evaluated using the same balanced dataset as specified in the Application Domains section.
We evaluate the accuracy of ENBED in generating mutations, using an encoder-decoder Transformer with Byte-Pair Encoding (BPE) tokenization (used in previous genomic models Fishman et al. (2023)) as a baseline. We compare against BPE because this method shares similarities with byte-level tokenization by starting with the basic {A, C, T, G} alphabet, but tries to optimize the vocabulary by combining simpler words into more complex ones based on the corpus the tokenizer is trained on. The training corpus itself is identical to the one used in pre-training ENBED, with the only difference being the tokenization procedure. While this procedure reduces the average number of tokens generated from any input sequence, it also results in reduced accuracy since modifying even a single base pair will output a significantly different tokenized sequence.
Top-1 and Top-5 Accuracy (%) scores are calculated by comparing predictions with real-world data from the Influenza Virus Resource Bao et al. (2008), with any deviation from an exact match being classified as incorrect. Top-5 scores are calculated by selecting the best candidate from the procedure described in Sec 2.5. Additionally, we also train a version of ENBED with the encoder removed, as a comparison of the sequence-to-sequence task performance between decoder-only and encoder-decoder models.
4 Ablation Studies
4.1 Encoder-decoder architecture
5 Discussion
The ENBED model demonstrates significant improvements over existing approaches in several areas of genomic sequence analysis. The encoder-decoder architecture, combined with byte-level tokenization and high-quality pre-training data, contributes to enhanced performance across multiple tasks. ENBED’s performance on established benchmarks is noteworthy, surpassing state-of-the-art results in 21 out of 25 tasks across the Nucleotide Transformer and Genomic Benchmarks datasets. This broad improvement suggests that our approach captures underlying genomic patterns more effectively than previous models. Additionally, the model successfully identified sequences containing noise with an accuracy of 97.6%, demonstrating its sensitivity to small-scale genomic perturbations. This is likely due to the byte-level tokenization approach used in ENBED, which allowed for accurate detection of variations at single-nucleotide resolutions.
The encoder-decoder structure proves particularly effective for sequence-to-sequence tasks like mutation generation. Our results show that ENBED outperforms baseline models in predicting Influenza virus mutations, achieving a top-5 accuracy of 95.4%. This was a significant improvement over the baseline model using byte-pair encoding (BPE) tokenization (56.1%), and another variant of ENBED without the encoder (72.1%). We chose to vary both the tokenization scheme and architecture in these cases while keeping the rest of the design choices unchanged in order to isolate the impact of these two factors. We find that the choice of BPE tokenization significantly impacts the model’s ability to generate mutations accurately, with byte-level tokenization providing a clear advantage due to its ability to capture single-nucleotide changes. We also see that an encoder-decoder architecture is crucial for this task, as the decoder-only model does not perform as well, following the trend observed in other sequence-to-sequence tasks Raffel et al. (2020); Fu et al. (2023).
It is also worth noting that the use of higher-quality pre-training data, including telomere-to-telomere assemblies, may contribute to ENBED’s improved performance. This comprehensive genomic representation likely allows the model to learn from previously underrepresented genomic regions. A study of the Nucleotide Transformer benchmarks (Table 12) with two versions of ENBED trained on different reference assemblies (GRCh38 and T2T-CHM13) showed that the model trained on the higher-quality T2T-CHM13 assembly outperformed the GRCh38 model across the board. This suggests that the choice of reference assembly can significantly impact the model’s performance, and that the use of more complete and accurate reference genomes can lead to better generalization.
Future work on this model could explore additional applications in genomics, such as variant effect prediction and protein structure studies.
6 Data and Code Availability
(GCF_000001635.27) and P. vivax (GCF_000002415.2) were obtained from NCBI RefSeq O’Leary et al. (2016) in FASTA format. Variant Calling Files (VCFs) for the 1000 Genomes Project Consortium (2015) were obtained from the European Bioinformatics Institute. Gene annotations were obtained from GENCODE Harrow et. al (2012) and Ensembl Martin et. al. (2022). The mutation tree was derived from the data assembled by the authors of Berman et al. (2020), sourced from the NCBI’s Influenza Virus Resource Bao et al. (2008).
The source code used to develop and fine-tune the foundation model has been released on Github  [^1] https://github.itap.purdue.edu/Clan-labs/ENBED and the weights of the model used in evaluation are available here  [^2] https://huggingface.co/malusare
7 Supplementary Material
The supplementary material (below) contains additional details on the model architecture, data sources, training procedures and evaluation metrics.
8 Acknowledgements
This work was supported in part by the National Science Foundation under grant [FW-HTF-R-2129097]; and the Anvil supercomputer Song et al. (2022) at Purdue University through allocation CIS230228 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants [2138259, 2138286, 2138307, 2137603, and 2138296]. The authors gratefully acknowledge the Walther Cancer Foundation and support from the Purdue University Institute for Cancer Research, [P30CA023168].
NT Benchmark Enformer DNABERT-2
NT
HyenaDNA
(1 Kb)
@@ 8 Acknowledgements
H3 0.719 0.785 0.791 0.779 0.723 0.802
H3K14ac 0.288 0.516 0.537 0.612 0.537 0.636
H3K36me3 0.344 0.591 0.616 0.613 0.611 0.624
H3K4me1 0.291 0.511 0.544 0.512 0.498 0.591
H3K4me2 0.211 0.336 0.322 0.455 0.433 0.501
H3K4me3 0.158 0.352 0.408 0.549 0.580 0.587
H3K79me3 0.496 0.613 0.621 0.672 0.648 0.756
H3K9ac 0.420 0.542 0.550 0.581 0.427 0.590
H4 0.732 0.796 0.807 0.763 0.750 0.823
H4ac 0.273 0.463 0.489 0.564 0.548 0.605
Promotor (all) 0.909 0.943 0.950 0.920 0.906 0.961
Promotor (non-TATA) 0.909 0.944 0.952 0.921 0.892 0.959
Promotor (TATA) 0.920 0.910 0.919 0.882 0.883 0.944
Splice acceptor 0.829 0.950 0.973 0.915 0.754 0.943
Splice donor 0.814 0.926 0.974 0.898 0.835 0.911
Enhancer 0.451 0.516 0.548 0.517 0.577 0.585
Enhancer Types 0.309 0.423 0.450 0.386 0.459 0.482
Table 1: Nucleotide Transformer (NT) Benchmarks. We evaluate our model using the 10-fold mean Matthews Correlation Coefficient (MCC) of the best performing variants of the Enformer Avsec et al. (2021), DNABERT Zhou et al. (2023), Nucleotide Transformer v2 Dalla-Torre et al. (2023), and HyenaDNA Nguyen et al. (2023), highlighting the best and second-best scores. The scores are sourced from a leaderboard maintained by the authors of Dalla-Torre et al. (2023) on the Hugging Face platform InstaDeepAI (2023).
Genomic Benchmark CNN DNABERT GPT
HyenaDNA
Nguyen et al. (2023)
@@
Mouse Enhancers 69.0 66.9 80.1 85.1 81.1 90.3
Human Enhancers (Cohn) 69.5 74.0 70.5 74.2 70.8 71.2
Human Enhancers (Ensembl) 68.9 85.7 83.5 89.2 90.2 92.2
Coding vs Intergenomic 87.6 92.5 88.8 91.3 90.7 93.0
Human vs Worm 93.0 96.5 95.6 96.6 94.4 97.3
Human Regulatory Elements 93.3 88.1 91.5 93.8 85.6 90.2
Human Promoter (Non-TATA) 84.6 85.6 87.7 96.6 90.4 97.2
Human OCR (Ensembl) 68.0 75.1 73.0 80.9 76.2 81.9
Table 2: Genomic Benchmarks. Accuracy (%) scores of the best and second-best model in the Genomic Benchmarks datasets Grevsova et al. (2022). The baseline CNN and GPT scores was calculated by the authors of Grevsova et al. (2022) and Nguyen et al. (2023) respectively.
@@
DNABERT Ji et al. (2021) 84.9
Nucleotide Transformer Dalla-Torre et al. (2023) 91.8
ENBED This paper 97.6
@@
DNABERT Ji et al. (2021) 63.2
Nucleotide Transformer Dalla-Torre et al. (2023) 67.5
HyenaDNA Nguyen et al. (2023) 72.8